Find the resultant of the following two vectors: i) 50 units due east and ii) 100 units 30 north of west.
1
Expert's answer
2021-09-28T11:11:45-0400

First, find the components of the resultant:


"R_E=50-100\cos30°=-71.6,\\
R_N=0+100\sin30°=50.\\
R=\sqrt{R_E^2+R_N^2}=87.3.\\\space\\
\theta=\arctan\frac{R_N}{R_E}=34.9°\text{ North of West.}"